OpenStudy (anonymous):

how do i find the limit as x approaches pi/2 of (sinx/x)

OpenStudy (anonymous):

\[\lim_{x\to \frac{\pi}{2}} \frac{\sin x}{x}\] What is the value that sin(x) assumes when x approaches pi/2 ? What is the value that x assumes when x approaches pi/2? (trivial) The limit of a ratio is the ratio of the limits \[\lim_{x \to \frac{\pi}{2}} \frac{\sin x}{x} = \frac{\lim_{x \to \frac{\pi}{2}} \sin x}{\lim_{x \to \frac{\pi}{2}} x}\] Can you complete the exercise?

OpenStudy (waynex):

That function is continuous everywhere except at x=0. That means this limit is calculated by plugging pi/2 into the function to get the answer directly.

OpenStudy (anonymous):

1/(pi/2) = 2/pi